I can shovel a driveway in 45 minutes; my brother can shovel the driveway in 36 minutes. If we work together, how long will it take?
My answer: 44.4 minutes Is this correct?

Answer by stanbon(75887)

I can shovel a driveway in 45 minutes; my brother can shovel the driveway in 36 minutes. If we work together, how long will it take?
My answer: 44.4 minutes Is this correct?
----
Your rate: 1/45 job/min
Bro's rate: 1/36 job/min
---
Together rate: 1/x job/min
-----
Equation:
rate + rate = together rate
1/45 + 1/36 = 1/x
-----
36x + 45x = 36*45
81x = 36*45
x = 4*5
x = 20 minutes (time to do the job together))
===================
Cheers,
Stan H.